40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Jul 22 20:51:13 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1862.JPGIMG_1863.JPG
    Reply  Tue Jul 26 18:42:16 2011, raphael, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1866.JPGIMG_1867.JPG
       Reply  Thu Jul 28 01:54:21 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts 2011_07_28.pngIMG_1869.JPGIMG_1874.JPGIMG_1875.JPG
          Reply  Fri Jul 29 00:42:19 2011, frank, tara, DailyProgress, optic, cleaning opto mechanical parts IMG_1880.JPGIMG_1883.JPGIMG_1889.JPGIMG_1886.JPGIMG_1878.JPG
             Reply  Tue Aug 2 23:58:40 2011, tara, DailyProgress, optic, New RCAV setup is locked IMG_1911.JPGrcav_2011_08_02.jpg
                Reply  Tue Aug 9 00:39:30 2011, tara, DailyProgress, optic, RCAV modematching optimized 
                   Reply  Mon Aug 15 23:13:53 2011, tara, DailyProgress, optic, RCAV modematching optimized 2011_08_15.png
                      Reply  Wed Aug 17 20:41:13 2011, tara, DailyProgress, optic, ACAV path is up layout_2011_08_17.jpg
Message ID: 644     Entry time: Thu Jul 28 01:54:21 2011     In reply to: 640     Reply to this: 647
Author: frank, tara 
Type: DailyProgress 
Category: optic 
Subject: cleaning opto mechanical parts 

  Today we started working on the layout. There is one mistake in the layout, the mirror behind AOM for REFCAV is too close to the insulation box, so we have to fix the layout.

 

 The oil on the table actually comes from holes on the table. About 5-6 screw holes had lot of oil, so I flushed them with methanol a few times.

IMG_1869.JPG

The HEPA filter is plugged in and turned on. I unplugged one of the monitors and used the outlet for the filter.

 

 The window is measured to be ~ 6 inches in diameter. Thus, the assumption in the design that the centers between two cavities are 3 inches is ok. If necessary, it can go up to 4 or 5 inches.

IMG_1875.JPG

IMG_1874.JPG

The layout is updated. The spot size in both AOMs are adjusted to 220 um.

2011_07_28.png

ELOG V3.1.3-